OpenStudy (anonymous):

Find the y-intercept for this linear equation. 7y + 3x - 21 = 0

OpenStudy (anonymous):

y=-3/7x+3 y=-3/7(0) + 3 y=3 y-intercept = (0,3)

OpenStudy (anonymous):

y intercept put x = 0 solve \[7y-21=0\] \[y=21\] \[y=3\]